How many numbers are possible from the set {0, 1, 2, 3, 4, 5, 6, 7} if:
a) the numbers must be less than 500?
b) the numbers must be odd?
c) the numbers must be even?

Answer by Edwin McCravy(20054)

You didn't specify whether we may repeat digits or not.
I will assume that we cannot.  If we can then post again
stating that digits may be repeated.

Assuming digits cannot be repeated:


a) the numbers must be less than 500?
Counting 1 and 2 digit numbers;
We consider 1 digit numbers as 2-digit numbers with 
left-most digit 0.

Choose units digit 8 ways {0,1,2,3,4,5,6,7}
Choose tens digit 7 ways

That's 8*7 = 56  1 and 2 digit numbers, counting 0.

Counting three digit numbers:
Choose hundreds digit 4 ways {1,2,3,4}
Choose tens digit 7 ways.
Choose units digit 6 ways
That's 4*7*6 = 168

So that's 56+168 = 224 counting 0.

-------------------------

b) the numbers must be odd?
I suppose you mean it can have any number of digits up through 8 digits,
since you didn't state otherwise.

7 or 8 digit numbers;
(We consider 7 digit numbers as 8 digit numbers whose left-most
digit is 0.)

Choose units digit 4 ways {1,3,5,7}
Choose tens-digit 7 ways
Choose hundreds digit 6 ways
Choose thousands digit 5 ways
Choose ten thousands digit 4 ways
Choose hundred thousands digit 3 ways
Choose millions digit 2 ways
Choose ten millions digit 1 way

That's 4*7*6*5*4*3*2*1 
------------
6 or 5 digit numbers
(We consider 5 digit numbers as 6 digit numbers whose left-most
digit is 0.)

Choose units digit 4 ways {1,3,5,7}
Choose tens-digit 7 ways
Choose hundreds digit 6 ways
Choose thousands digit 5 ways
Choose ten thousands digit 4 ways
Choose hundred thousands digit 3 ways

That's 4*7*6*5*4*3

----------------------  
Similarly, the number of 3 or 4 digit numbers is 4*7*6*5
----------------------
Similarly, the number of 1 or 2 digit numbers is 4*7 = 28
----------------------
Total: 4*7*6*5*4*3*2*1 + 4*7*6*5*4*3 + 4*7*6*5 + 4*7 = 31108
 
-------------------
c) the numbers must be even?

It's the same as (b) since there are same number even digits
{0,2,4,6) and there are odd digits {1,3,5,7} for the units
digit. Also 31108.
Edwin